# Out of spuriousity: Improving robustness to spurious correlations without group annotations

### v1

@@ 6.5 Ablation study
Without fine-tuning (setting $(4)$) the performance drops significantly. Without fine-tuning the model on the dataset $\mathcal{D}_{task}$, the subnetwork’s performance is similar to that before pruning. Note, that we only need to finetune $2-5$ epochs. This finding matches with previous work on modular subnetworks (Csordás et al., 2020; Zhang et al., 2021; Park et al., 2023) that require retraining or fine-tuning the remaining network after pruning to achieve better performance.
Subnetwork.
The three ablation experiments (settings $(5)$ to $(7)$) answer the question of whether one can improve the worst-group accuracy only by fine-tuning and getting advantages from the design of a class-balanced subset. We either fine-tune the whole network (setting $(5)$ and $(6)$), or only one last layer (setting $(7)$) for same number of epochs. The latter setting is equivalent to DFRclass, with $\mathcal{D}_{task}$. We observe no significant improvement when fine-tuning the whole network compared to fine-tuning only the last layer (setting 6 vs. 7), fine-tuning the whole network even results in worse WGA. As noted before, using the contrastive loss during fine-tuning again drops the group-worst accuracy.
Table 5: Ablation study on CelebA dataset (target: blond hair, spurious attribute: male). Effect of pruning (Pr) and finetuning (Ft) steps with (✓) and without (✗) our contrastive loss. ✓last means finetuning on the last layer only.
Setting Pruning Fine-tuning ConLoss Pr ConLoss Ft WGA AVG UA
1 (PruSC) 89.6 90.1 90.5
2 67.7 83.7 80.4
3 41.8 91.5 73.9
4 43.7 89.6 74.4
5 67.4 88.8 81.6
6 78.9 87.9 87.2
7 (DFRclass) last 79.5 83.8 85.5
6.6 Variant of contrastive learning.
@@
We study the effectiveness of our choices for contrastive batch learning in PruSC. Results reported in Tab. 6 shows that other ablations of contrastive batch are less effective than our choice in terms of improvement worst-group accuracy. Recall our default setting: sampling a random anchor, positives are defined by samples taking from different clusters and having the same class label with the anchor, while negatives are defined by samples from the same cluster with the anchor.
@@
Table 6: Average accuracy (AVG) and worst-group accuracy (WGA) of models trained on CelebA dataset with different choices of contrastive batch.
@@
Neg. Ablation 90.2% 73.1%
SupCon 92.2% 61.5%
We ablate the negative choice by keeping anchor and positives the same as default while negatives are chosen by the same cluster and different class label (Neg. Ablation). The effectiveness of this variant is affected by the purity of clusters. As shown in Fig. 2 and Tab. 1, $k-$means clustering on train representation of a convergent model can lead to high purity clusters in terms of class labels; therefore, by choosing different class negatives, we either lacking negative samples for contrastive learning or an early-stopping models is required to get more negative samples which we want to avoid. We also compare our choice to the classic SupCon (Khosla et al., 2020), i.e., positives and negatives are defined only by class labels. SupCon does not work since it aims to tighten samples within one cluster which has a high potential of having the same spurious features. Furthermore, SupCon only helps to separate classes, which ERM models in some cases do already well enough (cf. Fig. 2, left and Fig. 5, left). Our contrastive batch is defined not only by class labels like in SupCon but also by the positions of samples in the representation space based on clustering.
6.7 The impact of pruning ratio.
Figure 4: Worst-group accuracy (WGA) and average accuracy (AVG) of CelebA according to pruning ratio.
Fig. 4 shows the average and worst-group accuracy evaluated on the CelebA dataset with respect to different pruning ratios. We consider the set up of classifying blond hair and spurious feature is male or not-male. While the average accuracy does not change much as long as the pruning ratio is under $70\%$, the worst-group accuracy varies more significantly. The smaller gap between the average and worst-group accuracy indicates that the models are less prone to spurious correlations. In the CelebA dataset, the most effective pruning ratio is around $40\%$ to $50\%$. When pruning beyond $70\%$ of the parameters, the models seem to fail to predict accurately, as both the worst-group and average accuracy drop.
7 Feature visualization
@@
In this section, we discuss how PruSC affects the representation space at the end. We show in the previous part that ERM tends to learn and cluster samples based on spurious attributes during training (Fig. 2). In Fig. 5 (left), we visualize the embedding space of the non-blond hair class in the test set of the same ERM model, which reveals again ERM learns based on spurious features. Despite achieving high average test accuracy (90%), the fact that ERM learns to use spurious features for predicting leads to low worst-group accuracy ($49.7\%$) and lack of generalization ability.
@@
DCWP mitigates spurious correlations by extracting a subnetwork by training a contrastive loss, and sampling positives and negatives from bias-conflict (ERM misclassified) and bias-align (ERM correctly classified) cases. From the predictive result, DCWP successfully improves the worst-group accuracy ($73\%$), however, the model still tends to group samples by female attribute (Fig. 5, center). We hypothesize that defining bias-conflict based on ERM misclassified cases is not reliable. When a spurious attribute (female) is easy to learn and strongly correlated with a class (non-blond hair), the model easily learns and predicts correctly based on the spurious feature, even in the early stage of training.
@@
PruSC eliminates this reliance on spurious feature learning. While average accuracy remains high ($91\%$) through cross-entropy loss, the use of contrastive learning within clusters makes the learning spurious process harder, thus, improving the generalization. Our pruned model mixes up samples with different attributes within one class showing its independence on these spurious attributes (Fig. 5, right).
In our visualization, we use the official train and test split of CelebA where multiple spurious correlations are possible but not exposed with high spuriousness. We expect the phenomenon can be shown more concretely in the case of synthetic data constructed to have high spurious correlations. More visualizations are shown in Appendix A.4.
@@
Figure 5: The test set embedding space of ERM (left), DCWP (center) and PruSC (right) on CelebA for predicting blond hair color. We visualize the class of non-blond hair with colors representing three spurious attributes: female, young, and beard. All models achieve high average test accuracy (over 88%, see Tab. 2). However, while ERM and DCWP cluster samples by spurious attributes, our PruSC mixes samples with these attributes, resulting in better worst-group accuracy.
@@ 8 Conclusion
In this paper, we presented PruSC to extract a spurious-free subnetwork from a fully trained neural network. We do not require annotations of spurious features, but instead, optimize the representation space to uncorrelated spurious attributes and class labels by forcing examples of the same class which is in different clusters (presumably due to different spurious attributes) to become more similar. PruSC achieves competitive worst-group accuracy across several benchmarks without requiring annotations of the spurious feature(s). We find that this not only reduces annotation overhead and makes the method applicable in cases where spurious attributes are not known beforehand, but also results in models that are robust to multiple spurious correlations.
Notably, we extract the subnetwork from a fully trained network without extensively retraining the weights. This shows that even when models are heavily influenced by spurious correlations, there exists a smaller part of the network that learns the invariant features relevant to the task. This finding is a promising direction for future research.
One limitation of our pruning method is that it fails to explicitly explain the exact roles of the pruned connections. It is possible that the pruned connections encode some information essential for the classification task but not defined during pruning. This potential missing information could be crucial for effectively learning the WaterBird dataset, as discussed in Sec.6.3.
@@ A.1 Datasets
WaterBird.
WaterBird is an artificial dataset generated based on the code from (Sagawa et al., 2020a). Similar to the CelebA dataset, the training set of WaterBird suffers from both high class imbalance and significant spurious correlations, with waterbirds typically appearing in water backgrounds and landbirds in land backgrounds. Interestingly, the validation and test sets of WaterBird are balanced in terms of these spurious correlations, with equal numbers of waterbirds and landbirds appearing in both land and water backgrounds. Therefore, we expect that all methods tuning hyperparameters on the validation set will have access to distribution information reflective of the test set, leading to high overall performance and improved worst-group accuracy. We show details number of samples in each group in Tab. 8.
Table 8: Distribution of samples in dataset WaterBird and ISIC. Gray color highlights the minority group. WaterBird has similar distribution between validation and test sets, and they are both spurious-free within a class (i.e., equal number of samples with land or water background). In ISIC, we construct a group balanced validation set for the purpose of training DFR which requires a held-out group-balanced dataset.
@@
WaterBird waterbird on water 1057 133 642
@@ A.2 Task-oriented Subnetworks
Csordás et al. (2020) aims to highlight sets of non-shared weights solely responsible for individual class $k$ by training a mask with class $k$ removed from the training dataset ($\mathcal{C}_{task}=\mathcal{D}\setminus\{(x,y|y=k\}$) and need not add any other loss term. The weights solely responsible for this class will be absent from the resulting mask. Results show that the performance of the target class drops significantly, while only a small drop in performance is observed for non-target classes. This indicates a heavy reliance on class-exclusive, non-shared weights in the feature detectors. Interestingly, Csordás et al. (2020) also shows the mis-classification behaviors based on "shared features" (or possibly spurious features), e.g., when "airplane" is removed, images are often misclassified as "birds" or "ships," both of which commonly feature a blue background.
Zhang et al. (2021) investigates the capability of finding a subnetwork for a specific task within a trained dense network by extracting it using a specially designed held-out dataset tailored for that task. Specifically, the authors train a dense network on the Colored MNIST dataset, which contains a high spurious correlation between digits and colors, with each digit being strongly correlated with a particular color. Using the same technique as Csordás et al. (2020), but while pruning, Zhang et al. (2021) trains the network with a dataset balanced in terms of colors and digits to isolate subnetworks that solely predict either digits or colors. Surprisingly, even though the trained dense network suffers significantly from spurious correlations, it is still possible to extract a subnetwork that predicts only the digits with high accuracy, without extensive retraining on a balanced (or non-spurious) dataset.
Park et al. (2023) constructs the mask training data by upweighting the complex cases (called by bias-conflict samples) defined by the misclassified cases of the ERM model, aiming to extract a set of weights more robust to those cases therefore, eliminating bias in the network.
@@ A.3 Multiple spurious attributes
We analyze 7 chosen attributes having annotations from dataset, namely: male, no beard, heavy makeup, wearing lipstick, young, pale skin, big nose. Treating each attribute independently, we show their potential of being spurious attribute by comparing the worst-group accuracy from set set $\mathcal{G}$ = {(blond, $a_{i}$), (non-blond, $a_{i}$), (blond, non-$a_{i}$), (non-blond, non-$a_{i}$)} for each attribute $a_{i}$ and the ERM average accuracy. A fair model should treat all groups equally and minimize the influence of distribution. As shown in Tab. 9 (cf. ERM), we see some gaps between worst-group and average accuracy, and at the same time the unequal accuracy among groups.
@@
Tab. 9 shows more details on the accuracies of each group and compares among more methods; namely ERM, DCWP, GroupDRO, DFR, and Ours. Interestingly, when considering the group having the worst performance by ERM, our method outperforms all others by considerable margins. We argue that our cluster sampling stage can find hard cases even more effectively than methods using ERM failures such as DCWP.
Table 9: Details of group accuracy (%) with respect to multiple attributes in CelebA dataset. Average accuracy of baseline ResNet18 is 89.4%. After pruning, our subnetwork has keep_ratio 45.5% of ResNet18 weights. We color the column corresponding to the minority group with respect to the combination of target and spurious attribute.
T = Blond_hair {T=0, A=0} {T=0, A=1} {T=1, A=0} {T=1, A=1}

### v2

@@ 7.1 Component Ablation
Without fine-tuning after subnetwork extraction (setting $4$) the performance of the subnetwork drops significantly, to similar performance as before pruning. This finding aligns with previous work on modular subnetworks (Csordás et al., 2020; Zhang et al., 2021; Park et al., 2023), which require retraining or fine-tuning the extracted network after pruning to achieve better performance.
Contrastive Loss during Pruning.
Settings 2 and 3 show the effect of the contrastive loss during pruning (ConLoss Pr), which is the key feature of PruSC. In setting 3, we prune the networks using the cross-entropy loss and sparse regularization, i.e., vanilla pruning. Vanilla pruning leads to slightly better AVG but keeps the WGA the same as before pruning, while PruSC (setting 2) significantly increases the WGA, showing that the contrastive loss helps to find the correct subnetwork that is robust to spurious correlations. The contrastive loss forces the pruning process to retain only those connections that contribute to an optimized representation, ensuring that all spurious attributes are distributed in the representation space.
Contrastive Loss during Fine-tuning.
Fine-tuning with contrastive loss (ConLoss Ft) reduces the WGA significantly compared to fine-tuning without contrastive loss, both with (setting 1 vs. 2) and without subnetwork extraction (5 vs. 6 and 7). In all those settings, we fine-tune with the balanced subset $\mathcal{D}_{\text{task}}$. Kirichenko et al. (2023) showed that ERM models can effectively learn all features, but spurious features contribute more strongly to the prediction, due to the data distribution. Hence, during fine-tuning without subnetwork extraction, the model only has to adjust the classifier weights to rely more on invariant features. This is also confirmed by the increased WGA in setting 7 (fine-tuning only the last layer vs. fine-tuning all layers in setting 6). Adding the contrastive loss (setting 5), the model now has to optimize a dual objective: classification and representation, which is likely confusing rather than helpful, in particular as the balanced subset is comparably small. Similarly, adding the contrastive loss during fine-tuning after the extraction of a sub-network (setting 1) hurts performance. In this setting, the representation space has been optimized by the extraction of the subnetwork already and fine-tuning is only supposed to adjust the weight magnitudes after pruning. We hypothesize that forcing the network to also re-learn the representation space does not effectively align the remaining weights, but rather distorts weights and representations.
Contrastive Learning with (pseudo) Group Labels.
Instead of using contrastive learning with clusters defined by representation clustering, we can use annotated group labels. The clusters in $\mathcal{D}_{\text{task}}$ by ground truth group labels correspond to spurious attributes, i.e. the two clusters in CelebA with respect to the spurious attribute gender. This approach does significantly improve the WGA (from 50.2% to 84.8%, see Tab. 10 in the Appendix), highlighting the potential of using annotated or pseudo-group labels. However, contrastive learning with ground truth labels does not outperform PruSC with representation clustering (WGA 89.6%). We hypothesize that this is due to the alignment between our contrastive loss and contrastive batch sampling, which is particularly beneficial in the representation space. Further results and discussion are presented in App. A.5.
7.2 Impact of the Number of Clusters
Figure 4: Analysis of PruSC wrt. the number of clusters and pruning ratio on the CelebA dataset. (a) Average, minimum and maximum worst group accuracy (WGA) among 5 runs across varying numbers of clusters. (b) WGA and average accuracy (AVG) of varying pruning ratios.
We analyzed the impact of the number of clusters $k$ in $k$-means clustering on the CelebA dataset. Fig. 4(a) shows the average WGA together with minimum and maximum among 5 runs across different choices of $k$.
For $k<8$ performance is unstable and generally low. The effectiveness of PruSC depends on the purity of clusters with respect to classes and spurious features. For lower values of $k$, both, class purity ($\mathcal{Y}$-purity) and spurious purity ($\mathcal{A}$-purity) are also lower (cf. $k=4$ vs $k=8$, Tab. 1).
With lower cluster purity, it is more likely that each cluster contains samples from all groups in $\mathcal{G}$, making it harder for our method to sample a more group-balanced subset, which in turn affects WGA performance. For $k\geq 8$, the performance of PruSC remains rather stable, indicating that $k$ is not a critical hyper-parameter of our method, as long as it is sufficiently large to obtain pure clusters, both in terms of classes and spurious features.
7.3 Impact of the Pruning Ratio
Fig. 4(b) shows the average and worst-group accuracy of our model on CelebA for different pruning ratios. We test on the standard setup of classifying hair color and considering gender as spurious attribute. While the average accuracy does not change much as long as the pruning ratio is below $70\%$, the worst-group accuracy varies more significantly. Subnetworks in the regime of a small gap between average and worst-group accuracy are less prone to spurious correlations. On the CelebA dataset, the most effective pruning ratio is around $40\%$ to $50\%$. When pruning beyond $70\%$ of the parameters, the model seems to fail to predict accurately, as both, worst-group and average accuracy drop.
7.4 Choice of Positive and Negative Samples in Contrastive Learning
@@
We study the effectiveness of our sampling choices for contrastive batch learning in PruSC. Recall our default setting: for a random anchor, we sample positives from different clusters having the same class label as the anchor and negatives from the same cluster (regardless of class label). We compare this choice with sampling negatives from the same cluster, but different class labels (Negative Ablation) and classic supervised contrastive learning (SupCon) (Khosla et al., 2020), where positives and negatives are defined only by class labels. The results in Tab. 6 show that both choices are less effective w.r.t. worst-group accuracy.
@@
Table 6: Average accuracy (AVG) and worst-group accuracy (WGA) of PruSC with different sampling choices for the contrastive batch.
@@
Negative Ablation 90.2% 73.1%
SupCon (Khosla et al., 2020) 92.2% 61.5%
Sampling negatives from the same cluster as the anchor but with a different class label (Negative Ablation) is affected by the purity of clusters. As shown in Fig. 2 and Tab. 1, $k$-means clustering on the representations of a converged ERM model leads to high purity clusters in terms of class labels. Therefore, a cluster may not contain a sufficient number of negatives with different class labels. Early stopping could solve this problem, but requires access to the model training procedure, and requires additional optimization of the stopping criteria to balance model performance and the amount of misclassifications.
Instead, SupCon aims to bring the representations of the same class closer together, tightening the samples within the same cluster, which has a high potential to increase the spurious correlation instead of reducing it. Furthermore, SupCon primarily helps to separate classes (cf. increased average accuracy), which ERM models may already do sufficiently well (cf. Fig. 2, left and Fig. 5, left).
Our contrastive batch is defined not only by class labels as in SupCon, but also by the positions of the samples in the representation space based on clustering.
7.5 Representation Space after Mitigating Spurious Correlations
@@
In this section, we discuss how the mitigation strategy of DCWP and PruSC affects the representation space. In Sec. 5.1 we showed that ERM tends to learn and cluster samples based on spurious attributes during training (Fig. 2). In Fig. 5 (left), we visualize the embedding space of the non-blond hair class in the test set of the same ERM model, which reveals again ERM learns based on spurious attributes. Despite achieving high average test accuracy (90%), the fact that ERM relies on spurious correlations in its predictions leads to a low worst-group accuracy ($49.7\%$) and lack of generalization ability.
@@
DCWP mitigates spurious correlations by extracting a subnetwork by training a contrastive loss, and sampling positives and negatives from bias-conflict (early-stop ERM misclassified) and bias-align (early-stop ERM correctly classified) cases. In terms of performance measures, DCWP successfully improves worst-group accuracy ($73\%$), but the model still tends to group samples by the female attribute (Fig. 5, center). We hypothesize that defining bias-conflict based on ERM misclassified cases is not reliable. When a spurious attribute (female) is easy to learn and strongly correlated with a class (blond hair), the model learns easily and predicts correctly based on the spurious attribute, even in the early stages of training.
@@
PruSC eliminates this reliance on spurious feature learning. While the average accuracy remains high ($91\%$) through cross-entropy loss, the use of contrastive learning with a distortion of spurious clusters in the representation space eliminates the reliance on spurious features and thus improves generalization.
Our pruned model mixes up samples with different attributes within one class showing its independence from these spurious attributes (Fig. 5, right). Additional visualizations are shown in Appendix A.4.
@@
Figure 5: Test set representation space on CelebA for predicting hair color. The visualization shows data points of the class ‘non-blond hair‘, colors represent the spurious attributes. All models achieve high average test accuracy (over 88%, see Tab. 2). While ERM (left) and DCWP (center) cluster samples by spurious attributes, PruSC (right) mixes samples with these attributes, obtaining a better worst-group accuracy.
@@ 8 Conclusion
In this paper, we introduced PruSC (Pruning Spurious Correlations) to extract a spurious-free subnetwork from a fully trained dense neural network. Our method does not rely on annotations of spurious attributes, but instead optimizes the representation space to produce class-induced clusters rather than clusters induced by spurious attributes (as observed in the representation space of the dense network trained with ERM).
PruSC achieves competitive worst-group accuracy across several benchmarks without requiring any annotation of spurious features at all (not even in the validation set). While PruSC outperforms all annotation-free methods in comparison, we found that the spurious-free validation set of the Waterbirds dataset might lead to an overestimation of the performance of methods that rely on annotations in the validation set. Our truly annotation-free approach not only eliminates the annotation overhead and makes PruSC applicable to implicit and unknown spurious features, but also results in subnetworks that are robust to multiple spurious correlations.
By extracting the subnetwork from a fully trained network without extensive retraining of the weights, we show that even dense networks, that are heavily influenced by spurious correlations, contain smaller parts that learn invariant attributes relevant to the task. In future work, we plan to analyze the extracted subnetworks in detail, in order to quantify the information (encoded features and relations) that is retained and also the information that is lost. In particular, we aim to identify whether relevant information is lost and how to retain it (and vice versa for irrelevant information) in order to further improve performance.
@@ A.1 Datasets
Waterbirds.
Waterbirds is an artificial dataset generated based on the code from (Sagawa et al., 2020a). Similar to the CelebA dataset, the training set of Waterbirds suffers from both high class imbalance and significant spurious correlations, with waterbirds typically appearing in water backgrounds and landbirds in land backgrounds. Interestingly, the validation and test sets of Waterbirds are balanced in terms of these spurious correlations, with equal numbers of waterbirds and landbirds appearing in both land and water backgrounds. Therefore, we expect that all methods tuning hyper-parameters on the validation set will have access to distribution information reflective of the test set, leading to high overall performance and improved worst-group accuracy. We show details number of samples in each group in Tab. 8.
Table 8: Distribution of samples in dataset Waterbirds and ISIC. Gray color highlights the minority group. Waterbirds has similar distribution between validation and test sets, and they are both spurious-free within a class (i.e., equal number of samples with land or water background). In ISIC, we construct a group balanced validation set for the purpose of training DFR which requires a held-out group-balanced dataset.
@@
Waterbirds waterbird on water 1057 133 642
@@ A.2 Task-oriented Subnetworks
Csordás et al. (2020) aims to highlight sets of non-shared weights solely responsible for individual class $k$ by training a mask with class $k$ removed from the training dataset ($\mathcal{D}_{\text{task}}$ $\mathcal{D}\setminus\{(x,y|y=k\}$) and need not add any other loss term. The weights solely responsible for this class will be absent from the resulting mask. Results show that the performance of the target class drops significantly, while only a small drop in performance is observed for non-target classes. This indicates a heavy reliance on class-exclusive, non-shared weights in the feature detectors. Interestingly, Csordás et al. (2020) also shows the misclassification behaviors based on "shared features" (or possibly spurious features), e.g., when "airplane" is removed, images are often misclassified as "birds" or "ships," both of which commonly feature a blue background.
Zhang et al. (2021) investigate the capability of finding a subnetwork for a specific task within a trained dense network by extracting it using a specially designed held-out dataset tailored for that task. Specifically, the authors train a dense network on the Colored MNIST dataset, which contains a strong spurious correlation between digits and colors, with each digit being strongly correlated with a particular color. Using the same technique as Csordás et al. (2020), but while pruning, Zhang et al. (2021) train the network with a dataset balanced in terms of colors and digits to isolate subnetworks that solely predict either digits or colors. Surprisingly, even though the trained dense network suffers significantly from spurious correlations, it is still possible to extract a subnetwork that predicts only the digits with high accuracy without extensive retraining on a balanced (or non-spurious) dataset.
Park et al. (2023) construct the mask training data by upweighting the complex cases (called by bias-conflict samples) defined by the misclassified cases of the ERM model, aiming to extract a set of weights more robust to those cases, therefore eliminating bias in the network.
@@ A.3 Multiple spurious attributes
We analyze 7 chosen attributes having annotations from dataset, namely: male, no beard, heavy makeup, wearing lipstick, young, pale skin, big nose. Treating each attribute independently, we show their potential of being spurious attribute by comparing the worst-group accuracy from set set $\mathcal{G}$ = {(blond, $a_{i}$), (non-blond, $a_{i}$), (blond, non-$a_{i}$), (non-blond, non-$a_{i}$)} for each attribute $a_{i}$ and the ERM average accuracy. A fair model should treat all groups equally and minimize the influence of distribution. As shown in Tab. 9 (cf. ERM), we see some gaps between worst-group and average accuracy, and at the same time, the unequal accuracy among groups.
@@
Tab. 9 shows more details on the accuracies of each group and compares among more methods, namely ERM, DCWP, GroupDRO, DFR, and PruSC. Interestingly, when considering the group having the worst performance by ERM, our method outperforms all others by considerable margins. We argue that our cluster sampling stage can find hard cases even more effectively than methods using ERM failures such as DCWP.
Table 9: Details of group accuracy (%) with respect to multiple attributes in CelebA dataset. Average accuracy of baseline ResNet18 is 89.4%. After pruning, our subnetwork has keep-ratio 45.5% of ResNet18 weights. We color the column corresponding to the minority group with respect to the combination of target and spurious attribute.
T = blond hair {T=0, A=0} {T=0, A=1} {T=1, A=0} {T=1, A=1}